An infinite sheet of charge has a certain charge density. (A) If a charge of 7.60 nC experiences a force of 225 N from the shWhat am I doing wrong? Thank you.
Answer #1

E=F/Q

=225/7.6*10^-9

=29.6*10^9 N/C---a

b)E=sigma/2e0

sigma=29.6*10^9*2*8.85*10^-12

=0.52392 C/m^2---b